# Transport through a lattice with local loss: from quantum dots to lattice gases

### v1

@@ V.3 Extended transient effects near transmission resonances
Figure 5: Zero-temperature fermion flow over time in finite systems with different $N$ as in Fig. 4, for $\mu_{L}$ near the first conductivity step at $\mu_{L}=-\sqrt{2}$, and times less than $t=2.5N$. Top row: $N=20$ sites in the lattice ends, with initial Fermion numbers 3,4,5,6 from left plot to right. Middle row: $N=40$, initial numbers 8,9,10,11. Bottom row: $N=40$, initial numbers $19,20,21,22$. The different initial particle numbers all correspond to $\mu_{L}$ near $-\sqrt{2}$; in particular the second column of plots have $\mu_{L}$ just below $-\sqrt{2}$ and the third column have $\mu_{L}$ just above it. Since the group velocities at frequencies near these $\mu$ are near $\sqrt{2}$ rather than 2 as at $\omega=0$, finite-size effects are not yet apparent. Finite-time effects, however, are clear: the initial transient epoch lasts much longer near the conductivity resonance.
@@ V.4 Quasi-stationary current versus $\mu_{L}$
While finite lattices may thus force finite-time problems upon us, though, these problems may not actually be too severe. Fig. 6 shows finite-system analogs of Fig. 2a), with three different cases of $N$ (20, 40, and 80 in Figs. 6a), 6b), and 6c), respectively). Each of these three plots further shows three different simple attempts to extract the quasi-steady current from the growth of $\Delta\mathcal{N}(t)$ before the onset of finite-size effects at $t=N/\tau$. These attempts represent $I$ as local slopes $[\Delta\mathcal{N}(t_{m}+1/\tau)-\Delta\mathcal{N}(t_{m}-1/\tau)]/(2\tau)$, for three different choices of the measurement time $t_{m}$: the early choice $t_{m}=N/(4\tau)$, the late choice $t_{m}=3N/(4\tau)$, and the intermediate choice $t_{m}=N/(2\tau)$. If $\Delta\mathcal{N}(t)$ were as straight as it is in Fig. 4, at $\mu_{L}=0$, these three choices of $t_{m}$ would all yield very similar $I$ estimates. Discrepancies between the three attempts to find $I$ show the non-stationary nature of the fermion flow at other $\mu_{L}$.
@@
Figure 6: Zero-temperature current $I$ versus initial left-side chemical potential $\mu_{L}=\mu$ computed at three different times before the onset of finite-size effects, for each of three different finite lattice sizes. All cases have $\tau_{1}=0.1$, $\gamma=0.5$, $M=3$, and initial $\mu_{R}<-2$ (no particles in the right reservoir section). All three plots a), b), c) show the same thick, gray curve, which is $I(\mu)$ for the steady state of the infinite lattice with $\Lambda=1$ and $V=0$. In plots a), b), and c) we show cases with $N=$ 20, 40, 80. Each plot shows $I$ computed as $\big{(}[n_{R}(t+1)-n_{L}(t+1)]-[n_{R}(t-1)-n_{L}(t-1)]/2$ for $t=0.25N$ (o markers), $t=0.5N$ (+ markers), and $t=0.75N$ (x markers). The markers are discrete points rather than continuous lines because zero-temperature fermions in a finite lattice simply occupy the $n$ lowest modes, for integer $n$ between 1 and $N$.
@@
As Fig. 6 shows, it is possible to resolve the steady-state current $I(\mu_{L})$ from Fig. 2a) fairly well with finite lattices as small as $N=20$, and lattices as large as $N=80$ allow essentially duplicating the steady-state current curve of the infinite system, with clear conductivity steps that are noticeably smoother for the even-parity channels that are affected by loss. Even though extended transient effects near the steps still exist even for $N=80$, as shown in Fig. 5, this issue only really affects the two data points on either side of the step, and for $N=80$ there are many other points to map the curve out accurately.
@@
Ultimately the main limitation from finite lattices turns out to be the fundamentally limited frequency resolution from only having an $N$-dimensional single-particle Hilbert space in an $N$-site lattice. As long as $N$ is larger than $M$ by enough both to delay finite-size effects until the initial transient epoch is mostly over and to resolve the $12M+$ 1 conductivity steps in the frequency band $-2<\omega<2$, experiments with ultracold Fermi gases in finite optical lattices can indeed reveal the same interesting transport phenomena with a lossy site that have been predicted for finite chains of quantum dots coupled to leads.
@@ VI Discussion and Outlook
In conclusion, we have constructed a class of fermionic lattice gas models which can reproduce the results of [2] in the limit $\Lambda\to\infty$, and then represent quantum gas experiments in the limit $\Lambda\to 1$. Different but similar effects are predicted for quantum gases in optical lattice, in comparison with the parity-dependent conductivity steps that are predicted to be seen in quantum dot chains between leads. Finite-temperature effects seem likely to be a practical challenge in near-future experiments with quantum gases. While these effects will not prevent observation of the most interesting conductivity features, they will likely prevent quantum gas experiments from matching the sharpness of conductivity steps that can be seen in condensed matter systems.
One new possibility that can appear with quantum gases in lattices is transport of bosons instead of fermions. In the absence of interactions, however, bosons do not readily display the dramatic step-like conductivity of zero-temperature fermions, because Bose-Einstein distributions cannot replicate the moveable sharp edge of a low-temperature Fermi-Dirac distribution with variable Fermi energy. For non-interacting bosons, the chemical potential can rise no higher than the bottom of the single-particle energy band; the only way to have varying exposure to the different discrete transport channels through the finite middle chain is to vary the reservoir temperatures.
Figure 7: Current $I$ through the finite chain of seven sites ($M=3$) as in Fig. 2, but for non-interacting bosons at maximal chemical potential $\mu_{L}=-2$, as a function of temperature. No conductivity steps are apparent: while the sharp edge of the zero-temperature Fermi-Dirac distribution rises with increasing $\mu$, increasing temperature in the Bose-Einstein distribution merely spreads the distribution out flatter, steadily overlapping more and more with all seven conductivity resonances.
As Fig. 7 disappointingly shows, increasing temperature simply mixes in higher-frequency channels continuously. No conductivity steps can be seen—at least not in scenarios like those we have considered in this paper, and not without interactions among the bosons. Perhaps different kinds of experiments will be able to show more abrupt unlocking of discrete transport channels for bosons. Further studies may also reveal more interesting transport behavior in interacting bosons, which can, among other complications, have higher chemical potentials.

### v2

@@ V.3 Extended transient effects near transmission resonances
Figure 6: Zero-temperature fermion flow over time in finite systems with different $N$ as in Fig. 5, for $\mu_{L}$ near the first conductance step at $\mu_{L}=-\sqrt{2}$, and times less than $t=2.5N$. Top row: $N=20$ sites in the lattice ends, with initial Fermion numbers 3,4,5,6 from left plot to right. Middle row: $N=40$, initial numbers 8,9,10,11. Bottom row: $N=80$, initial numbers $19,20,21,22$. The different initial particle numbers all correspond to $\mu_{L}$ near $-\sqrt{2}$; in particular the second column of plots have $\mu_{L}$ just below $-\sqrt{2}$ and the third column have $\mu_{L}$ just above it. Since the group velocities at frequencies near these $\mu$ are near $\sqrt{2}$ rather than 2 as at $\omega=0$, finite-size effects are not yet apparent. Finite-time effects, however, are clear: the initial transient epoch lasts much longer near the conductance resonance.
@@ V.4 Quasi-stationary current versus $\mu_{L}$
While finite lattices may thus force finite-time problems upon us, though, these problems may not actually be too severe. Fig. 7 shows finite-system analogs of Fig. 3a), with three different cases of $N$ (20, 40, and 80 in Figs. 7a), 7b), and 7c), respectively). Each of these three plots further shows three different simple attempts to extract the quasi-steady current from the growth of $\Delta\mathcal{N}(t)$ before the onset of finite-size effects at $t=N/\tau$. These attempts represent $I$ as local slopes $[\Delta\mathcal{N}(t_{m}+1/\tau)-\Delta\mathcal{N}(t_{m}-1/\tau)]/(2\tau)$, for three different choices of the measurement time $t_{m}$: the early choice $t_{m}=N/(4\tau)$, the late choice $t_{m}=3N/(4\tau)$, and the intermediate choice $t_{m}=N/(2\tau)$. If $\Delta\mathcal{N}(t)$ were as straight as it is in Fig. 5, at $\mu_{L}=0$, these three choices of $t_{m}$ would all yield very similar $I$ estimates. Discrepancies between the three attempts to find $I$ show the non-stationary nature of the fermion flow at other $\mu_{L}$.
@@
Figure 7: Zero-temperature current $I$ versus initial left-side chemical potential $\mu_{L}=\mu$ computed at three different times before the onset of finite-size effects, for each of three different finite lattice sizes. All cases have $\tau_{1}=0.1$, $\gamma=0.5$, $M=3$, and initial $\mu_{R}<-2$ (no particles in the right reservoir section). All three plots a), b), c) show the same thick, gray curve, which is $I(\mu)$ for the steady state of the infinite lattice with $\Lambda=1$ and $V=0$. In plots a), b), and c) we show cases with $N=$ 20, 40, 80. Each plot shows $I$ computed as $\big{(}[n_{R}(t+1)-n_{L}(t+1)]-[n_{R}(t-1)-n_{L}(t-1)]/2$ for $t=0.25N$ (o markers), $t=0.5N$ (+ markers), and $t=0.75N$ (x markers). The markers are discrete points rather than continuous lines, because fermions at zero temperature in a finite lattice simply occupy the $n$ lowest modes, for integer $n$ between 1 and $N$, and so the initial state only actually changes with chemical potential $\mu$ in discrete steps.
@@
As Fig. 7 shows, it is possible to resolve the steady-state current $I(\mu_{L})$ from Fig. 3a) fairly well with finite lattices as small as $N=20$, and lattices as large as $N=80$ allow essentially duplicating the steady-state current curve of the infinite system, with clear conductance steps that are noticeably smoother for the even-parity channels that are affected by loss. Even though extended transient effects near the steps still exist even for $N=80$, as shown in Fig. 6, this issue only really affects the two data points on either side of the step, and for $N=80$ there are many other points to map the curve out accurately.
@@
Ultimately the main limitation from finite lattices turns out to be the fundamentally limited frequency resolution from only having an $N$-dimensional single-particle Hilbert space in an $N$-site lattice. As long as $N$ is larger than $M$ by enough both to delay finite-size effects until the initial transient epoch is mostly over and to resolve the $12M+$ 1 conductance steps in the frequency band $-2<\omega<2$, experiments with ultracold Fermi gases in finite optical lattices can indeed reveal the same interesting transport phenomena with a lossy site that have been predicted for finite chains of quantum dots coupled to leads.
@@ VI Discussion and Outlook
In conclusion, we have constructed a class of fermionic lattice gas models which can reproduce the results of [10] in the limit $\Lambda\to\infty$, and then represent quantum gas experiments in the limit $\Lambda\to 1$. Different but similar effects are predicted for quantum gases in optical lattice, in comparison with the parity-dependent conductance steps that are predicted to be seen in quantum dot chains between leads. Finite-temperature effects seem likely to be a practical challenge in near-future experiments with quantum gases. While these effects will not prevent observation of the most interesting conductance features, they will likely prevent quantum gas experiments from matching the sharpness of conductance steps that can be seen in condensed matter systems.
As in the condensed matter context of [10], we have also focused in this paper on the symmetry dependence of loss effects, whereby localised loss smoothes out some conductance steps while leaving other steps sharp because their associated transport channel is protected from loss by symmetry. In particular we have considered loss at the central site only, so that channels with odd parity are unaffected. The flexibility of cold atom experiments may also open new avenues, however, to using tailored loss for more general control over quantum transport, such as by tuning loss at multiple sites in order to filter out more transport channels, or by making the loss rates time-dependent in order to select particular wave packets for transport while absorbing all others.
One entirely new possibility that can appear with quantum gases in lattices is transport of bosons instead of fermions. In the absence of interactions, however, bosons do not readily display the dramatic step-like conductance of zero-temperature fermions, because Bose-Einstein distributions cannot replicate the moveable sharp edge of a low-temperature Fermi-Dirac distribution with variable Fermi energy. For non-interacting bosons, the chemical potential can rise no higher than the bottom of the single-particle energy band; the only way to have varying exposure to the different discrete transport channels through the finite middle chain is to vary the reservoir temperatures.
Figure 8: Current $I$ through the finite chain of seven sites ($M=3$) as in Fig. 3, but for non-interacting bosons at maximal chemical potential $\mu_{L}=-2$, as a function of temperature. No conductance steps are apparent: while the sharp edge of the zero-temperature Fermi-Dirac distribution rises with increasing $\mu$, increasing temperature in the Bose-Einstein distribution merely spreads the distribution out flatter, steadily overlapping more and more with all seven conductance resonances.
As Fig. 8 disappointingly shows, increasing temperature simply mixes in higher-frequency channels continuously. No conductance steps can be seen—at least not in scenarios like those we have considered in this paper, and not without interactions among the bosons. Perhaps different kinds of experiments will be able to show more abrupt unlocking of discrete transport channels for bosons. Further studies may also reveal more interesting transport behavior in interacting bosons, which can, among other complications, have higher chemical potentials.